4. Find the solution set for the following equation. Check for extraneous solutions:

Square root of 8 - 7x = -2(x + 1)

Answer by fractalier(6550)

From
sqrt(8 - 7x) = -2(x + 1)
Let us square both sides first
8 - 7x = 4(x + 1)^2
8 - 7x = 4(x^2 + 2x + 1)
8 - 7x = 4x^2 + 8x + 4
now collect like terms and solve
4x^2 + 15x - 4 = 0
(4x - 1)(x + 4) = 0
x = 1/4 or x = -4
You plug them in to the original equation to check.
x = 1/4 does not work but x = -4 does.